"You need $73323 in 11 years. If you can earn 0.25 percent per month, how much will you have to deposit today (in $)?"

The correct answer is $52735.36. But I keep getting $52970.10.

My calculations are:

PMT=0

FV= $73323

I/Y= 0.25 x 12= 3

N= 11

I am hoping you can explain to me where my error is.
Answer #1

When you do I/Y=0.25*12=3% you are assuming effective annual rate is 3%..Instead it will be (1+0.25%)^12-1=3.0416%..So Instead of 3 use 3.0416% and use N=11

Alternatively, use N=11*12 and I/Y=0.25%